OpenStudy (anonymous):

how do i solve... the square root of 3x - 5= x - 5

OpenStudy (lgbasallote):

\[\sqrt{3x - 5} = x -5\] ???

OpenStudy (barrycarter):

Assuming @lgbasallote said it correctly, start by squaring both sides

OpenStudy (anonymous):

yes... how did you get the square root sign

OpenStudy (lgbasallote):

by following your instructions?

OpenStudy (anonymous):

ok whatever...

OpenStudy (anonymous):

ok i squared both sides

OpenStudy (anonymous):

hello??

OpenStudy (barrycarter):

What did you get after squaring?

OpenStudy (anonymous):

3x-5=(x+-5)(x+-5

OpenStudy (barrycarter):

OK, and when you multiply it out, that's a quadratic equation, right?

OpenStudy (anonymous):

yeah i think... i got x^2+-13x+30... is that right?

OpenStudy (barrycarter):

Yes, it does. (you can always check final answers at wolframalpha.com)

OpenStudy (anonymous):

so is that all i do or do i solve for x...???

OpenStudy (barrycarter):

You have to solve by x either by factoring, completing the square, or applying the quadratic formula

OpenStudy (anonymous):

i am trying to make it to where someone could do foil...but im having difficulties factoring it in...!

OpenStudy (barrycarter):

What two numbers multiply out to 30, but add to -13 (hint: -3 is one of the numbers)

OpenStudy (anonymous):

o is it -3 and -10

OpenStudy (anonymous):

so it would be (x-3)(x-10)?????

OpenStudy (barrycarter):

Correct, (x-3)*(x-10) = 0

OpenStudy (anonymous):

so is that all my answer is when it tells me to solve.

OpenStudy (gw2011):

x^2-13x+30 = 0 factoring the left side you get: (x-10 (x-3) = 0 x-10 = 0 x = 10 x-3 = 0 x = 3 When you substitute each of these values,x=10 and x=3, into the original equation, then the only one that has the right solution is; x=10 For x=10 you get from the original equation 5=5 But for x=3 you get from the original equation 2=-2 which is not correct

OpenStudy (anonymous):

o ok thank you everyone!!!! :)

OpenStudy (barrycarter):

Good catch, @gw2011 -- I forgot about extraneous roots for a moment